A Company placed into service a forklift in 2018 which costs $20,000; assume 7 year life. Using MACRS show the depreciation for 2018, 2019 and 2020. Show the adjusted basis after the 2020 depreciation.
Answer #1

Calculation of Depriciation

2018 = $20000 / 7 = $2857

2019 = ($20000 - $2857) / 6 = $2857

2020 = ($20000 - $5714) / 6 = $2857